(1 point) A 7 year $16000 par-valued bond pays semi-annual coupons. If the yield rate is y2 is $11882.17, what is the coupon rate c2? 10% and the purchase price Answer:
Answer #1
Par value of Bonds 16000
Semi annual periods (7*2) 14
Market rate of interest 10%
Annuity PVF for 14 periods at 5% 9.89864
PVF at 14th period at 5% 0.505068
Let, Semi annual cash interest be x
Present value of maturity value 8081.088
Present value of Interest payments 9.89864 x
Price 8081.09 + 9.89864 x
As per given in question, issue price is $ 11882.17
Therefore, equation is as under:
11882.17 = 8081.09 + 9.89864 x
9.89864x = 3801.08
x =384
Semi annual cash interest = 384
Annualised Interest = 384 *2 = 768
Coupon rate of Interest = 768 /16000*100 = 4.80%